# Counterfactual Sensitivity in Quantitative Spatial Models

### v2

@@ Abstract
Counterfactuals in quantitative spatial models are functions of the current state of the world and the model parameters. Current practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, which is both practical and theoretically justified. I apply the proposed method to the applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022) and find non-trivial uncertainty about counterfactuals.
@@ 1 Introduction
Economists use quantitative spatial models to answer counterfactual questions. For example, what is the effect on welfare when trade costs between a set of countries decrease because of an infrastructure investment or a trade agreement? How will expenditure shares change if some countries obtain a novel technology which increases their production efficiency? These counterfactual questions generally have the same structure: what will happen to the endogenous variables when the exogenous variables change in some specific way?
@@
These quantitative spatial models are assumed to exactly match the data, and allow us to write the counterfactual change in an endogenous variable of interest as a function solely of the observables. However, the observables are often noisily measured, which causes meaningful uncertainty about counterfactual predictions. Furthermore, this generates a novel measurement error problem because the object of interest is a function of the true realized values of the observables rather than a function of the correctly measured distribution of the observables, as traditionally considered in the econometrics literature on measurement error.
@@
In this paper, to provide tools for quantifying uncertainty in this novel measurement error setting, I propose an empirical Bayes (EB) approach. This approach takes as inputs a model for the measurement error and a prior on the true data and allows researchers to incorporate economic knowledge through the prior. I outline how to calibrate the parameters of the prior using the data and find the posterior distribution for the counterfactual change in an endogenous variable of interest. From this posterior we can obtain credible sets. In the appendix I also develop an adversarial alternative to the EB approach, which derives worst-case bounds for the counterfactual under researcher-selected bounds on the measurement error.
@@
To fix ideas, consider as an example the canonical Armington model (Armington, 1969). Under this model, we can write the proportional change in welfare as result of a proportional shock to trade costs as a function of baseline bilateral trade flows, baseline incomes and the trade elasticity. We do not know the trade elasticity and hence estimate it using the observed baseline bilateral trade flows. The question I will aim to answer is how mismeasurement in the bilateral trade flows generates uncertainty around the predicted change in welfare.
As illustrated by the Armington model, the structural parameters of the model are often estimated using the noisily measured data. Hence, in the presence of measurement error we must account for estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure. To address this, we aim to cover the “oracle” confidence interval that accounts for estimation error, based on the correctly-measured data, with prescribed probability.
@@
The EB approach requires researchers to specify a model for the measurement error and a prior on the true data. I provide a widely applicable default approach, which proposes default choices for the measurement error and the prior on the true data when our object of interest is a function of a set of noisily measured positive flows between locations. Specifically, I recommend a normal-normal model where the prior is centered on a gravity model. This default approach can be applied out-of-the-box to most quantitative spatial models, but can also easily be adapted to specific settings.
@@
To illustrate the impact of accounting for measurement error, I revisit the applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For the counterfactual analysis in Adao, Costinot, and Donaldson (2017), which considers welfare effects from China joining the WTO, I model measurement error in the bilateral trade flows. I use the widely applicable default approach and calibrate the parameters of the prior and the measurement error model using the mirror trade dataset of Linsi, Burgoon, and Mügge (2023). This dataset collects the reported bilateral trade flows by both the exporter and importer, which I interpret as two noisy measures of the true bilateral trade flow. I plot the change in China’s welfare from 1996 to 2011, and add credible sets that account for measurement error and estimation error. These sets are wide, and for multiple time periods it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
@@
For the application from Allen and Arkolakis (2022), the counterfactual question of interest is which highway links in the United States have the highest return on investment, and are hence most worth improving. I model mismeasurement in traffic flows, and again use the widely applicable default approach. I calibrate the measurement error model based on estimates from Musunuru and Porter (2019). I calculate the credible sets that account for measurement error and estimation error for the three links with the highest return on investment. These sets are large, and when estimation error is considered we lose the ability to clearly rank the links.
@@
This paper contributes to the small literature on improving counterfactual calculations in quantitative spatial economics (Ossa, 2015; Adao, Costinot, and Donaldson, 2017; Adão, Costinot, and Donaldson, 2023). The most relevant paper in this literature is Dingel and Tintelnot (2020), which studies calibration procedures in granular settings. It considers models which in addition to assuming that the data are perfectly explainable, also assume there is a continuum of agents. Then, in the case where there are only a handful of observations, individual idiosyncrasies do not wash out and are incorporated in the model, which causes overfitting and poor performance out-of-sample. I focus on the complementary question of uncertainty quantification due to measurement and estimation error, which arises even in non-granular settings.
@@
I make a novel connection between the econometrics literature of measurement error modeling and the literatures on quantitative spatial economics. As is argued in Goes (2023) and Linsi, Burgoon, and Mügge (2023), measurement error is prevalent in the data underlying quantitative spatial models. The most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Schennach, 2016). The key difference between the current setting and conventional measurement error settings is that the object of interest directly depends on the correctly measured data observations rather than on the correctly measured distribution of the data. I discuss this further in Section 3.1.1 below.
@@
The rest of this paper is organized as follows. The next section introduces the notation and the setting I consider. Section 3 discusses measurement error in quantitative spatial models and introduces my empirical Bayes method for uncertainty quantification. Section 4 applies my method to the application in Adao, Costinot, and Donaldson (2017). Section 5 considers uncertainty quantification in the economic geography model from Allen and Arkolakis (2022). Section 6 concludes.
2 Counterfactuals in Quantitative Spatial Models
@@
This section introduces the notation I will be using and discusses the assumptions that commonly underly counterfactual analyses in quantitative spatial models.
@@ 2.1 Notation and Key Assumption
To state the key assumption on quantitative spatial models, I begin by discussing the setting without measurement error. Consider an equilibrium, denoted by $\left(X_{O},X_{U},N_{O},N_{U}\right)$. Here, $X_{O}\in\mathcal{X}_{O}\subseteq\mathbb{R}^{d_{X,O}}$ are exogenous observables, $X_{U}\in\mathcal{X}_{U}\subseteq\mathbb{R}^{d_{X,U}}$ are exogenous unobservables, $N_{O}\in\mathcal{N}_{O}\subseteq\mathbb{R}^{d_{N,O}}$ are endogenous observables, and $N_{U}\in\mathcal{N}_{U}\subseteq\mathbb{R}^{d_{N,U}}$ are endogenous unobservables. Exogenous variables are fundamentals that the model takes as given. Endogenous variables are the variables researchers try to explain, and are determined within the model. [^1] One might argue that “external” and “internal” are less ambiguous terms than “exogenous” and “endogenous”, respectively (LeRoy, 2004). I opt for the latter because of the convention in the literature on quantitative spatial models. These equilibrium variables are connected to each other through the equilibrium conditions. These are, for a given parameter $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$, given by
@@
Having observed a single draw $\left(X_{O},N_{O}\right)$ from the distribution $\mathcal{P}_{O}$, we are then interested in what happens to the endogenous variables $\left(N_{O},N_{U}\right)$ when we change the baseline exogenous variables $\left(X_{O},X_{U}\right)$ in a proportional way. Formally, for a given vector of exogenous change variables $\left(\hat{X}_{O},\hat{X}_{U}\right)$, we want to find a corresponding vector of endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$ such that
@@
where $\odot$ denotes element-wise multiplication. The ultimate object of interest, which I denote by $\hat{k}$, will then be some transformation of the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$, the observables $\left(X_{O},N_{O}\right)$ and the structural parameter $\theta$. It could for example be a specific endogenous change variable or a change in welfare. The key assumption that this counterfactual change variable of interest has to satisfy is:
@@
For a given counterfactual question of interest, we can write $\hat{k}$ as a function solely of the observables $\left(X_{O},N_{O}\right)$:
@@
Assumption 1 implies that if $\left(X_{O},N_{O}\right)$ is observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\hat{k}$. Loosely, in quantitative spatial models Assumption 1 often holds because, for a fixed counterfactual question of interest as described by $\left(\hat{X}_{O},\hat{X}_{U}\right)$, we can solve for the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$ as a function of the observables $\left(X_{O},N_{O}\right)$. The exact functional form of $g_{\hat{k}}$ depends on the specific quantitative spatial model that is considered. In Appendix A I discuss the relevant underlying assumptions for two leading classes of models, namely invertible models and exact hat algebra models.
@@ 2.1.1 Running Example: Armington Model
For illustration, I will now work through finding $g_{\hat{k}}$ in a simple example. The simplest workhorse model in trade economics is the Armington model (Armington, 1969), as for example outlined in Costinot and Rodríguez-Clare (2014). There are countries indexed by $i,j=1,...,n$, which make a distinct good with productivity $Q_{i}$. Preferences are CES, so utility, which equals real consumption, for country $j$ is given by
@@
These equations can be rearranged to find the function $f$ as in Equation (1). We can plug the second set of equations into the first set of equations and solve for income levels $\left\{Y_{i}\right\}$ as functions of the exogenous variables. By Walras’ Law, income levels are only pinned down up to a multiplicative constant. Having solved for $\left\{Y_{i}\right\}$, we can then also solve for $\left\{\lambda_{ij}\right\}$. The model exactly matches the data in that for given income levels $\left\{Y_{i}\right\}$ and expenditure shares $\left\{\lambda_{ij}\right\}$, there exist trade costs $\left\{\tau_{ij}\right\}$ and productivities $\left\{Q_{i}\right\}$ such that these equilibrium conditions hold exactly.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\hat{\tau}_{ij}\right\}$, holding productivities $\left\{Q_{i}\right\}$ constant. The unobserved counterfactual levels of the trade costs hence are $\left\{\tau_{ij}\hat{\tau}_{ij}\right\}$. In Appendix B.1 it is shown that we can find:
@@
By again plugging in the second set of equations into the first one, we can then solve for the changes in income levels $\left\{\hat{Y}_{i}\right\}$ which again, by Walras’ Law, are only pinned down up to a multiplicative constant, and subsequently solve for $\left\{\hat{\lambda}_{ij}\right\}$. Hence, for a fixed counterfactual question of interest $\left\{\hat{\tau}_{ij}\right\}$, we can solve for $\hat{N}_{O}=\left(\left\{\hat{Y}_{i}\right\},\left\{\hat{\lambda}_{ij}\right\}\right)$ as a function solely of the observables $N_{O}=\left(\left\{Y_{i}\right\},\left\{\lambda_{ij}\right\}\right)$ and the structural parameter $\theta=\varepsilon$.
@@
It follows that for each of these counterfactual change variables of interest, for a given counterfactual question as described by $\left\{\hat{\tau}_{ij}\right\}$, we only require knowledge of the observables $\left\{Y_{i}\right\}$ and $\left\{\lambda_{ij}\right\}$. For example, focusing on the change in welfare in the first country, if we assume that the trade elasticity $\varepsilon$ is known we find
@@
2.2 Incorporating Parameter Uncertainty
The counterfactual change variable of interest will generally depend on the structural parameter $\theta$. If we assume that this structural parameter is fixed and known, our object of interest is $\hat{k}=g_{\hat{k}}\left(X_{O},N_{O};\theta\right)$. In practice, however, we do not know the structural parameter exactly and we hence have to use an estimator $\tilde{\theta}$. This subsection shows that the notation from Assumption 1 is general, and can account for estimation error by finding the appropriate $g$-functions.
Suppose that, instead of the true parameter $\theta$, we use an estimator $\tilde{\theta}\left(X_{O},N_{O}\right)$. Our focus then shifts from the true counterfactual change variable of interest $\hat{k}$ to the point estimate
$\hat{k}^{E}=g_{\hat{k}^{E}}\left(X_{O},N_{O}\right)\equiv g_{\hat{k}}\left(X_{O},N_{O};\tilde{\theta}\left(X_{O},N_{O}\right)\right).$
If we know the sampling distribution of $\tilde{\theta}\left(X_{O},N_{O}\right)$, we can find a $100\left(1-\beta\right)\%$ frequentist confidence region around $\hat{k}^{E}$, which I will denote by $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$. Our object of interest then becomes this confidence region, and it follows that the relevant $g$-functions are
$L_{\hat{k}^{E}}^{\beta}$ $=g_{L_{\hat{k}^{E}}^{\beta}}\left(X_{O},N_{O}\right)\quad\mathrm{and}\quad U_{\hat{k}^{E}}^{\beta}=g_{U_{\hat{k}^{E}}^{\beta}}\left(X_{O},N_{O}\right).$

### v3

# Counterfactual Sensitivity in Quantitative Trade and Spatial Models

@@ Abstract
Counterfactuals in quantitative trade and spatial models are functions of the current state of the world and the model parameters. Common practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, and show that it is both practical and theoretically justified. I apply the proposed method to the settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022) and find non-trivial uncertainty about counterfactuals.
@@ 1 Introduction
Economists use quantitative trade and spatial models to answer counterfactual questions. For example, what is the effect on welfare when trade costs between a set of countries decrease because of an infrastructure investment or a trade agreement? How will expenditure shares change if some countries obtain a novel technology which increases their production efficiency? These counterfactual questions generally have the same structure: what will happen to endogenous variables when exogenous variables change in some specific way?
@@
Quantitative trade and spatial models exactly match the data, and allow us to write the counterfactual change in an endogenous variable of interest as a function solely of the observables. However, the observables are often noisily measured, which introduces uncertainty about counterfactual predictions. Furthermore, this generates a novel measurement error problem because the object of interest is a function of the true realized values of the observables, rather than a function of the correctly measured distribution of the observables as traditionally considered in the econometrics literature on measurement error.
@@
In this paper, I propose an empirical Bayes (EB) approach for quantifying uncertainty in this novel measurement error setting. This approach takes as inputs a model for the measurement error and a prior distribution on the true data, and allows researchers to incorporate economic knowledge through the prior. I outline how to calibrate the parameters of the prior distribution using the data and quantify uncertainty about the counterfactual change in an endogenous variable of interest.
@@
To fix ideas, consider as an example the canonical Armington model (Armington, 1969). Under this model, we can write the proportional change in welfare from a proportional shock to trade costs as a function of baseline bilateral trade flows, baseline incomes and the trade elasticity (Arkolakis, Costinot, and Rodríguez-Clare, 2012). We do not know the trade elasticity and hence estimate it using the observed baseline bilateral trade flows. The question I will aim to answer is how mismeasurement in the bilateral trade flows generates uncertainty around the predicted change in welfare. As illustrated by the Armington model, the structural parameters of the model are often estimated using the noisily measured data. Hence, we must account for estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure.
@@
The EB approach requires researchers to specify a measurement error model and a prior distribution on the true data. I provide a widely applicable default approach, which calibrates the measurement error model and the prior distribution based on observed data. Specifically, I recommend using a log-normal measurement error model and a log-normal prior which is centered on a gravity model. This default approach can be applied out-of-the-box to many quantitative trade and spatial models, but can also easily be adapted to other settings.
@@
To illustrate the impact of accounting for measurement error, I revisit the settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For the counterfactual analysis in Adao, Costinot, and Donaldson (2017), which considers welfare effects from China joining the WTO, I model measurement error in the bilateral trade flows. I use my default approach and calibrate the parameters of the prior distribution and the measurement error model using the mirror trade dataset of Linsi, Burgoon, and Mügge (2023). This dataset collects the reported bilateral trade flows by both the exporter and importer, which I interpret as two noisy measures of the true bilateral trade flow. I plot the change in China’s welfare from 1996 to 2011, and add intervals that account for measurement error and estimation error. These sets are wide, and for multiple time periods it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
@@
For the setting in Allen and Arkolakis (2022), the counterfactual question of interest is which highway links in the United States have the highest return on investment, and are hence most worth improving. I model mismeasurement in traffic flows, and again use my default approach, calibrating the measurement error model based on estimates from Musunuru and Porter (2019). I calculate the intervals that account for measurement error and estimation error for the three links with the highest return on investment. These sets are large, but the ranking of the three links is robust to measurement error and estimation error.
@@
This paper contributes to a small literature on improving counterfactual calculations in quantitative trade and spatial economics (Adao, Costinot, and Donaldson, 2017; Kehoe, Pujolas, and Rossbach, 2017; Adão, Costinot, and Donaldson, 2023). The most relevant paper in this literature is Dingel and Tintelnot (2020), which studies calibration procedures in granular settings. It considers models which, in addition to assuming that the data are perfectly explainable, also assume there is a continuum of agents. Then, in the case where there are only a handful of observations, individual idiosyncrasies do not wash out and are incorporated in the model, which causes overfitting and poor performance out-of-sample. I focus on the complementary question of uncertainty quantification due to measurement and estimation error, which arises even in non-granular settings.
@@
I make a novel connection between the econometrics literature on measurement error modeling and the literature on quantitative trade and spatial economics. As is argued in Goes (2023) and Linsi, Burgoon, and Mügge (2023), measurement error is prevalent in the data underlying quantitative trade and spatial models. The most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Schennach, 2016). The key difference between the current setting and conventional measurement error settings is that the object of interest directly depends on the correctly measured data observations rather than on the correctly measured distribution of the data.
@@
The rest of this paper is organized as follows. The next section introduces the notation and the setting I consider. Section 3 discusses estimation error and measurement error in quantitative trade and spatial models, discusses the relationship to the econometrics literature on measurement error, and introduces my empirical Bayes method for uncertainty quantification. Section 4 applies my method to the setting in Adao, Costinot, and Donaldson (2017). Section 5 considers uncertainty quantification in the economic geography model from Allen and Arkolakis (2022). Section 6 concludes.
2 Counterfactuals in Quantitative Trade and Spatial Models
@@
This section introduces the notation and discusses the assumptions that commonly underly counterfactual analyses in quantitative trade and spatial models.
@@ 2.1 Notation and Key Assumption
To state the key assumption in quantitative trade and spatial models, I begin by discussing the setting without estimation error or measurement error. Consider an equilibrium, denoted by $\left(X_{O},X_{U},N_{O},N_{U}\right)$. Here, $X_{O}\in\mathcal{X}_{O}\subseteq\mathbb{R}^{d_{X,O}}$ are exogenous observables, $X_{U}\in\mathcal{X}_{U}\subseteq\mathbb{R}^{d_{X,U}}$ are exogenous unobservables, $N_{O}\in\mathcal{N}_{O}\subseteq\mathbb{R}^{d_{N,O}}$ are endogenous observables, and $N_{U}\in\mathcal{N}_{U}\subseteq\mathbb{R}^{d_{N,U}}$ are endogenous unobservables. Exogenous variables are fundamentals that the model takes as given. Endogenous variables are the variables researchers try to explain, and are determined within the model. [^1] One might argue that “external” and “internal” are less ambiguous terms than “exogenous” and “endogenous”, respectively (LeRoy, 2004). I opt to use “exogenous” and “endogenous” because of the convention in the literature on quantitative trade and spatial models. These equilibrium variables are connected to each other through the equilibrium conditions. These are, for a given parameter $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$, given by
@@
Having observed a single draw $\left(X_{O},N_{O}\right)$ from the distribution $\mathcal{P}_{O}$, we are interested in what happens to the endogenous variables $\left(N_{O},N_{U}\right)$ when we change the baseline exogenous variables $\left(X_{O},X_{U}\right)$ in a proportional way. Formally, for a given vector of exogenous change variables $\left(\hat{X}_{O},\hat{X}_{U}\right)\in\mathcal{\hat{X}}_{O}\times\mathcal{\hat{X}}_{U}\subseteq\mathbb{R}^{d_{X,O}+d_{X,U}}$, we want to find a corresponding vector of endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)\in\mathcal{\hat{N}}_{O}\times\mathcal{\hat{N}}_{U}\subseteq\mathbb{R}^{d_{N,O}+d_{N,U}}$ such that
@@
where $\odot$ denotes element-wise multiplication.
The ultimate object of interest, which I denote by $\hat{k}$, will be some transformation of the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$, the observables $\left(X_{O},N_{O}\right)$ and the structural parameter $\theta$. It could for example be a specific endogenous change variable or a change in welfare. The key assumption that this counterfactual change variable of interest has to satisfy is:
@@
For a given counterfactual question of interest and known $\theta$, we can write $\hat{k}$ as a function solely of the observables $\left(X_{O},N_{O}\right)$:
@@
Assumption 1 implies that if $\left(X_{O},N_{O}\right)$ are observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\hat{k}$. Loosely, in quantitative trade and spatial models Assumption 1 often holds because, for a fixed counterfactual question of interest as described by $\left(\hat{X}_{O},\hat{X}_{U}\right)$, we can solve for the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$ as a function of the observables $\left(X_{O},N_{O}\right)$. The exact functional form of $g_{\hat{k}}$ depends on the specific quantitative model that is considered. In Appendix A I discuss the relevant sufficient conditions for Assumption 1 in two leading classes of models, namely invertible models and exact hat algebra models.
@@ 2.1.1 Running Example: Armington Model
For illustration, I will derive $g_{\hat{k}}$ in a simple example. The simplest workhorse model in trade economics is the Armington model (Armington, 1969), as for example outlined in Costinot and Rodríguez-Clare (2014). There are countries indexed by $i,j=1,...,n$, which are endowed with $Q_{i}$ units of a unique good. Preferences are CES, so utility, which equals real consumption, for country $j$ is given by
@@
These equations can be rearranged and stacked to find the function $f$ as in Equation (1). We can plug the second set of equations into the first set of equations and solve for income levels $\left\{Y_{i}\right\}$ as functions of the exogenous variables. By Walras’ Law, income levels are only pinned down up to a multiplicative constant. Having solved for $\left\{Y_{i}\right\}$, we can then also solve for $\left\{\lambda_{ij}\right\}$. The model exactly matches the data in that for given income levels $\left\{Y_{i}\right\}$ and expenditure shares $\left\{\lambda_{ij}\right\}$, there exist trade costs $\left\{\tau_{ij}\right\}$ and endowments $\left\{Q_{i}\right\}$ such that these equilibrium conditions hold exactly.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\hat{\tau}_{ij}\right\}$, holding endowments $\left\{Q_{i}\right\}$ constant. The unobserved counterfactual levels of the trade costs are hence $\left\{\tau_{ij}\hat{\tau}_{ij}\right\}$. In Appendix B.1 it is shown that we can find:
@@
By again plugging in the second set of equations into the first, we can solve for the changes in income levels $\left\{\hat{Y}_{i}\right\}$ which again, by Walras’ Law, are only pinned down up to a multiplicative constant, and subsequently solve for $\left\{\hat{\lambda}_{ij}\right\}$. Hence, for a fixed counterfactual question of interest $\left\{\hat{\tau}_{ij}\right\}$, we can solve for $\hat{N}_{O}=\left(\left\{\hat{Y}_{i}\right\},\left\{\hat{\lambda}_{ij}\right\}\right)$ as a function solely of the observables $N_{O}=\left(\left\{Y_{i}\right\},\left\{\lambda_{ij}\right\}\right)$ and the structural parameter $\theta=\varepsilon$.
@@
It follows that for each of these counterfactual change variables of interest, for a given counterfactual question as described by $\left\{\hat{\tau}_{ij}\right\}$, we only require knowledge of the observables $\left\{Y_{i}\right\}$ and $\left\{\lambda_{ij}\right\}$ and the trade elasticity $\varepsilon$. For example, focusing on the change in welfare in the first country, we have
@@
3 Empirical Bayes Uncertainty Quantification
This section introduces estimation error and measurement error to quantitative trade and spatial models. It proposes an empirical Bayes approach to uncertainty quantification, and outlines how to quantify uncertainty about the counterfactual change variable of interest.
3.1 Introducing Estimation Error